# Signal-to-noise Ratio Analytic Formulae of the Inspiral Binary Black Holes in TianQin

### v2

@@ Appendix A Detectors Setting
$S_{a}=$ $N_{a}\times\left[1+\left(\frac{0.4\times 10^{-3}~{}{\rm Hz}}{f}\right)^{2}\right]$
×$\times\left[1+\left(\frac{f}{8\times 10^{-3}~{}{\rm Hz}}\right)^{4}\right],$ (20)
@@
$t_{l}$ $=t_{c}-\frac{5}{(8\pi f_{\rm low})^{8/3}}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{-5/3},$
$\approx t_{c}-18~{}{\rm h}\times\left(\frac{10^{7}~{}M_{\odot}}{\mathcal{M}}\right)^{5/3}.$ (25)
$t_{h}$ $=t_{c}-\frac{5}{(8\pi f_{\rm high})^{8/3}}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{-5/3},$
$\approx t_{c}-180~{}{\rm h}\times\left(\frac{1~{}M_{\odot}}{\mathcal{M}}\right)^{5/3}.$ (26)
@@ Appendix B ASA SNR with low-frequency approximation
$\rho^{2}$ $=4\int^{f_{\rm max}}_{f_{\rm min}}{\frac{\mathcal{A}_{ASA}^{2}(f)}{S_{n}(f)}}{\rm d}f,$
$=\frac{\pi^{8/3}}{c^{3}}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}\left(G\mathcal{M}\right)^{5/3}\int^{f_{\rm max}}_{f_{\rm min}}f^{5/3}{\rm d}f,$
$=\frac{15c^{2}}{2048N_{a}t_{c}}\left(\frac{L}{D_{L}}\right)^{2}\frac{T_{\rm obs}}{t_{c}-T_{\rm obs}},$ (30)
@@
$\rho(T_{\rm obs})$ $=\sqrt{\frac{15}{2048}}\frac{L}{D_{L}}\frac{c}{\sqrt{N_{a}t_{c}}}\times\sqrt{\frac{t_{obs}}{t_{c}-t_{obs}}},$
$\approx 185.2\times\frac{1~{}{\rm Gpc}}{D_{L}}\sqrt{\frac{1~{}{\rm week}}{t_{c}}}\times\sqrt{\frac{T_{\rm obs}}{t_{c}-T_{\rm obs}}}.$ (31)
Appendix C Calculation of the response factor
@@
$G^{22}_{slr}(f,t)=$ $-\frac{{\rm i}\pi fL}{2c}\times{\rm sinc}\left[\frac{\pi fL}{c}(1-k\cdot n_{l})\right]$
×$\times\exp\left[{\rm i}\pi f\left(\frac{L+k\cdot(p_{r}+p_{s})}{c}\right)\right]\times P_{slr}(t),$ (33)
@@
By eliminating the phase terms that do not affect the SNR and employing a low-frequency approximation, the amplitude of the observable can be simplified as:
@@
where the $\tilde{a}$, $\tilde{e}$ and $\tilde{t}$ terms can be simplified, under the low-frequency approximation, to [40]:
@@
$S_{A}=$ $S_{E}$
$=$ $8\sin^{2}\left(\frac{f}{f_{\ast}}\right)\left[4\left(1+\cos\left(\frac{f}{f_{\ast}}\right)+\cos\left(2\frac{f}{f_{\ast}}\right)\right)S_{\rm acc}\right.$
$\left.+\left(2+\cos\left(\frac{f}{f_{\ast}}\right)\right)S_{\rm oms}\right],$ (42)
where $S_{\rm acc}=S_{a}\left(1/(2\pi fc)\right)^{2}$ is the acceleration noise, primarily dominant at low frequencies. And $S_{\rm oms}=S_{p}\left((2\pi f)/c\right)^{2}$ is the displacement or position noise, primarily dominant at high frequencies. Therefore, after also applying the low-frequency approximation and the frequency-independent $S_{a}$ approximation, the low-frequency TDI noise PSD of TianQin can be simplified to:
@@
$\rho^{2}$ $=4\int^{f_{\rm max}}_{f_{\rm min}}{\frac{\tilde{A}^{2}(f)+\tilde{E}^{2}(f)}{S_{A}}}{\rm d}f,$
$\simeq\frac{\pi^{8/3}}{2c^{3}}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}\left(G\mathcal{M}\right)^{5/3}\int^{f_{\rm max}}_{f_{\rm min}}P_{\rm res}^{2}f^{5/3}{\rm d}f,$ (44)
@@
It is noteworthy that, although each $P_{slr}$ fluctuates with time and frequency, the aggregation into $P_{\rm res}$ form renders it temporally invariant, with variation contingent only upon the sources’ sky positions and polar angles. This enables the extrication of this term from the integral, thereby maintaining the simplicity of the SNR formulation. Consequently, the SNR formula in the non-average form simply incorporates an extra $P_{\rm res}$ term compared to the ASA form (Equation III.1):

### v3

@@ 5 Conclusion
This paper considers a relatively ideal scenario. In the future, TianQin is likely to encounter issues such as data gaps. However, previous work[38] has shown that data gaps during the inspiral phase do not significantly change the SNR of BBHs. If necessary, adjusting the frequency integral boundaries for a piecewise calculation is straightforward. The total SNR can be represented as the square root of the sum of the squares of the SNRs of multiple data segments.
@@ Appendix A Detectors Setting
$S_{a}=$ $N_{a}\times\left[1+\left(\frac{0.4\times 10^{-3}~{}{\rm Hz}}{f}\right)^{2}\right]\times\left[1+\left(\frac{f}{8\times 10^{-3}~{}{\rm Hz}}\right)^{4}\right],$ (22)
@@
$t_{l}$ $=t_{c}-\frac{5}{(8\pi f_{\rm low})^{8/3}}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{-5/3}\approx t_{c}-18~{}{\rm h}\times\left(\frac{10^{7}~{}M_{\odot}}{\mathcal{M}}\right)^{5/3}.$ (27)
$t_{h}$ $=t_{c}-\frac{5}{(8\pi f_{\rm high})^{8/3}}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{-5/3}\approx t_{c}-180~{}{\rm h}\times\left(\frac{1~{}M_{\odot}}{\mathcal{M}}\right)^{5/3}.$ (28)
@@ Appendix B ASA SNR with low-frequency approximation
$\rho^{2}$ $=4\int^{f_{\rm max}}_{f_{\rm min}}{\frac{\mathcal{A}_{ASA}^{2}(f)}{S_{n}(f)}}{\rm d}f=\frac{\pi^{8/3}}{c^{3}}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}\left(G\mathcal{M}\right)^{5/3}\int^{f_{\rm max}}_{f_{\rm min}}f^{5/3}{\rm d}f,$ (32)
$=\frac{15c^{2}}{2048N_{a}t_{c}}\left(\frac{L}{D_{L}}\right)^{2}\frac{T_{\rm obs}}{t_{c}-T_{\rm obs}},$
@@
$\rho(T_{\rm obs})$ $=\sqrt{\frac{15}{2048}}\frac{L}{D_{L}}\frac{c}{\sqrt{N_{a}t_{c}}}\times\sqrt{\frac{t_{obs}}{t_{c}-t_{obs}}},$ (33)
$\approx 185.2\times\frac{1~{}{\rm Gpc}}{D_{L}}\sqrt{\frac{1~{}{\rm week}}{t_{c}}}\times\sqrt{\frac{T_{\rm obs}}{t_{c}-T_{\rm obs}}}.$
Appendix C Calculation of the correction
The error in our formulas within the low frequency region of MBHBs mainly comes from the frequency dependence of acceleration noise, as indicated in Equation 20 and 22. Considering the frequency dependence, the SNR squared becomes:
$\rho_{\rm corr}^{2}(t)\approx\frac{\mathcal{M}^{5/3}}{N_{a}\times D_{L}{}^{2}}\times\sum_{i=0}^{n}c_{i}\left[x(t)-x(0)\right]^{i-8/3},$ (34)
where $c_{i}$ represents distinct constants, and $x\equiv\mathcal{M}^{5/8}\left(t_{c}-t\right)^{3/8}$. It should be noted that for LISA, $c_{1}=0$, so its lowest-order correction is $i=2$.
Appendix D Calculation of the response factor
@@
$G^{22}_{slr}(f,t)=$ $-\frac{{\rm i}\pi fL}{2c}\times{\rm sinc}\left[\frac{\pi fL}{c}(1-k\cdot n_{l})\right]$ (36)
×$\times\exp\left[{\rm i}\pi f\left(\frac{L+k\cdot(p_{r}+p_{s})}{c}\right)\right]\times P_{slr}(t),$
@@
By eliminating the phase terms that do not affect the SNR and employing a LFA, the amplitude of the observable can be simplified as:
@@
where the $\tilde{a}$, $\tilde{e}$ and $\tilde{t}$ terms can be simplified, under the LFA, to [42]:
@@
$S_{A}$ $=$ $S_{E}$ (45)
$=$ $8\sin^{2}\left(\frac{f}{f_{\ast}}\right)\left[4\left(1+\cos\left(\frac{f}{f_{\ast}}\right)+\cos\left(\frac{2f}{f_{\ast}}\right)\right)S_{\rm acc}+\left(2+\cos\left(\frac{f}{f_{\ast}}\right)\right)S_{\rm oms}\right],$
where $S_{\rm acc}=S_{a}\left(1/(2\pi fc)\right)^{2}$ is the acceleration noise, primarily dominant at low frequencies. And $S_{\rm oms}=S_{p}\left((2\pi f)/c\right)^{2}$ is the displacement or position noise, primarily dominant at high frequencies. Therefore, after also applying the LFA and the frequency-independent $S_{a}$ approximation, the low-frequency TDI noise PSD of TianQin can be simplified to:
@@
$\rho^{2}$ $=4\int^{f_{\rm max}}_{f_{\rm min}}{\frac{\tilde{A}^{2}(f)+\tilde{E}^{2}(f)}{S_{A}}}{\rm d}f,$ (47)
$\simeq\frac{\pi^{8/3}}{2c^{3}}\left(\frac{L}{D_{L}}\right)^{2}\frac{1}{N_{a}}\left(G\mathcal{M}\right)^{5/3}\int^{f_{\rm max}}_{f_{\rm min}}P_{\rm res}^{2}f^{5/3}{\rm d}f,$
@@
It is noteworthy that, although each $P_{slr}$ fluctuates with time and frequency, the aggregation into $P_{\rm res}$ form renders it temporally invariant, with variation contingent only upon the sources’ sky positions and polar angles. This enables the extrication of this term from the integral, thereby maintaining the simplicity of the SNR formulation. Consequently, the SNR formula in the non-average form simply incorporates an extra $P_{\rm res}$ term compared to the ASA form (Equation 7):